Brad is nine more than two times as old as Ben. In 15 years the sum of their ages wil be 111. How old is Ben now?
Answer by laoman(51)

Let the age of brad be br and age of ben be b
From the first statement, br = 2*b + 9 ------ eqn 1
From the second statement, i.e in 15 yrs,
Brad's age in 15yrs is br+15 and Ben's age would be b+15
Given that the sum of their ages would be 111
i.e br+15+b+15 = 111
Simplifying
br+b = 111-30
br+b =81 --------eqn 2
Substituting 1 into 2,
2*b + 9 + b =81,
3b = 81-9 = 72
b = 72/3 = 24
Therefore Ben's age, b is 24 years
QED


Also we can find br, i.e br = 2*b+9 = 2*24+9 = 57
Therefore, Brad's age is also 57
In the next 15 years, their age sum would absolutely be 111